what are the equivalent fractions of the following:
1. 1/3
2.8/7
3.12/11
4.3/4
5.11/8
6.17/12

Answer by AnubhavRane(4)

There are infinitely many answeres possible. Multiply the numerator and denominator by the same number (except zero)
1. 1/3 = 11/33 ... multiplying both by 11
or 1/3 = 2/6 ... multiplying both by 2
got it?
So why not use the simplest number for multiplication ... TEN!
So,
2.8/7 = 80/70
3.12/11 = 120/110
4.3/4 = 30/40
5.11/8 = 110/80
6.17/12 = 170/120